# Geographic Variation in

### v1

@@ Abstract
Using complete-count register data spanning three generations, we compare inter- and multigenerational transmission processes across municipalities in Sweden. We first document spatial patterns in intergenerational (parent-child) mobility, and study whether those patterns are robust to the choice of mobility statistic and the quality of the underlying microdata. We then ask whether there exists similar geographic variation in multigenerational mobility. Interpreting those patterns through the lens of a latent factor model, we identify which features of the transmission process vary across places.
@@ 1 Introduction
An increasingly popular approach to study intergenerational mobility is to compare mobility rates across regions. Theoretical models with social or spatial dimensions (e.g. Durlauf, 1996 and Benabou, 1996) suggest that income persists across generations partly because parents shape their children’s social environment, such as schools or neighborhoods. Empirically, regional variations can be used to identify causal effects; for example, Holmlund (2008) and Pekkarinen, Uusitalo and Kerr (2009) exploit the staggered implementation of an educational reform to estimate its impact on intergenerational mobility. Following Chetty et al. (2014), some studies invert this research design: rather than focusing on a specific policy or event, they first map regional variation in intergenerational mobility and then search for regional correlates that could explain this variation. [^1] Examples include Connolly, Corak and Haeck (2019), Connolly, Haeck and Lapierre (2019) and Corak (2020) for Canada, Deutscher and Mazumder (2020) for Australia, Heidrich (2017) and Brandén (2019) for Sweden, Risa (2019) and Bütikofer, Dalla-Zuanna and Salvanes (2022) for Norway, Eriksen and Munk (2020) for Denmark, Acciari, Polo and Violante (2022) for Italy, or Bell, Blundell and Machin (2023) for the UK.
In this study, we extend this “regional approach” by considering not only inter- but also multigenerational associations. Such associations are interesting both from a descriptive and a causal perspective. Recent studies indicate that socioeconomic disparities are more persistent than extrapolations from the available parent-child correlations would suggest (Anderson, Sheppard and Monden, 2018; Solon, 2018; Stuhler, 2024). This contrast between inter- and multigenerational correlations is in turn informative about the statistical properties and potential mechanisms of mobility processes (e.g., Collado, Ortuño-Ortín and Stuhler, 2023). By adopting such multigenerational perspective, we aim to better understand the causes and implications of regional differences in mobility.
@@
We first report regional estimates of educational and earnings mobility across three generations for cohorts born in the 1980s across 290 municipalities. Municipalities are key geographical entities in the Swedish context, as many public goods and schooling are provided at this level, and often correspond to the relevant labor market of their inhabitants. Our evidence here is complementary to evidence provided by Heidrich (2017) and Brandén (2019), who estimate income mobility for earlier birth cohorts and a broader definition of local labor markets or commuting zones.
@@
We then study whether such regional rankings are sensitive to the choice of outcome (education vs. earnings), statistic (e.g., regression vs. correlation coefficients), time frame (inter- vs. multigenerational), or lineage (paternal vs. maternal line). One particular question is whether summary statistics based on education – which can be estimated in most settings – can serve as a substitute for more data-demanding measures based on income. This part of our analysis is complementary to recent work by Deutscher and Mazumder (2023), who compare the ranking of Australian regions across different measures of income mobility.
@@
We find that regional rankings based on inter- and multigenerational measures are similar. Municipalities with low intergenerational mobility in education (high parent-child correlations) also tend to have low multigenerational correlations (high grandparent-child correlations). In contrast, the association between education- and earnings-based mobility measures is weak. While this observation contrasts with cross-national comparisons, in which educational and income mobility are found to be closely related (Blanden, 2013; Stuhler, 2018), it is reminiscent of the sociological discourse on the “mobility paradox”, with some countries showing diverging patterns of income compared to educational or occupational mobility (e.g. Breen, Mood and Jonsson, 2016; Karlson, 2021).
@@
We next study whether multigenerational mobility and cross-sectional inequality are systematically related across regions. A well-documented pattern in international comparisons shows a negative association between income mobility and inequality (Blanden, 2013, Corak, 2013). We find that the same relationship holds across municipalities within Sweden, not only for inter- but also for multigenerational mobility. In line with evidence by Brandén (2019), who analyzed father-child mobility for earlier Swedish cohorts, we observe that municipalities that are characterized by high income inequality in the parent’s generation also tend to exhibit low mobility. However, the strength of this relationship varies considerably with the choice of mobility statistic. Educational mobility is more strongly correlated with inequality than income mobility, possibly due to greater measurement error in income data. Interestingly, the relationship with inequality is particularly pronounced for multigenerational correlations in education.
@@ 2.1 Samples and outcomes
Our measure of labor earnings includes the annual sum (before taxes and transfers) of all earnings according to employer-reported tax and earnings statements to the national tax agency, which captures all types of labor compensation subject to payroll taxes. [^5] We observe labor earnings (including income from self-employment) for all residents in 1968, 1970, 1971, 1973, 1975, 1976, 1979, 1980, 1982 and annually for 1985-2020. We adjust annual incomes for inflation using the CPI. While it would have been interesting to use disposable income, the information on disposable income does not go as far back in time. The measure is not censored or top-coded, and includes bonus payments. For our purposes, it is crucial to use earnings measures that are as comparable as possible over generations, as our estimates would be distorted if the influence of lifecycle and attenuation biases varies across generations. Given the structure of our data set, with earnings observed over a given time period and different generations (naturally) born several decades apart, it is not possible to measure earnings over the same age ranges for all generations. We therefore use a Mincer-type approach to predict prime-age earnings for all generations, which accounts for differences in income growth across education groups and differences in income levels across cohorts. [^6] While this approach improves our earnings measures, they are still unlikely to be fully comparable across generations. Specifically, we do not account for the observation that age-education-earnings profiles may differ across cohorts (Heckman, Lochner and Todd, 2006) or be steeper for children from parents with higher socioeconomic status (Mello, Nybom and Stuhler, 2025). However, we do allow for individual intercepts, thereby capturing general shifts in earnings levels across generations. Table A.1 in the Appendix compares area-level inter- and multigenerational earnings rank slopes for our prediction-based measure with more standard measures based on observed incomes. While the dispersion across municipalities is largely similar across measures, the prediction-based measure yields larger coefficients with magnitudes that are more in line with recent evidence for Sweden.
@@ 2.2 Mobility statistics
where $y_{ir}$ is one of the outcomes described above for individual $i$ in region (municipality) $r$ and $y_{i}^{p}$ is the corresponding outcome for a parent or grandparent. The slope $\beta_{r}$ measures “relative” mobility, comparing socioeconomic status relative to the status of others in the same generation. For education, we also consider the corresponding Pearson correlation coefficient, which abstracts from shifts in the variance between the two generations within the region. For our analysis of earnings mobility, we consider either log earnings (in which case the slope coefficient $\beta_{r}$ can be interpreted as the intergenerational elasticity, i.e. the IGE) or earnings ranks (in which case $\beta_{r}$ can be referred to as the rank slope). Ranks are defined on the national level, separately within each child cohort. [^10] Based on the estimated intercept $\hat{\alpha}_{r}$ and the slope $\hat{\beta}_{r}$ we further construct the expected “absolute upward mobility” of children born to parents at the 25th percentile (i.e., for $y_{i}^{p}=0.25$), as defined in Chetty et al. (2014). Our primary focus is on education correlations or the earnings rank slope, and we focus only occasionally on the IGE and the measure of upward mobility.
@@ 3 Geographical Variation in Mobility
As daughters’ education resembles their mother’s more than their father’s education (the relations are more symmetric for sons), mother-child correlations tend to be marginally larger than the corresponding father-child correlations (0.31 vs 0.30); see Appendix Table A.2 for separate statistics by child gender. The inverse pattern applies for grandparent correlations, which are systematically larger for grandfather-child than for grandmother-child pairs (0.12 vs 0.09). [^11] These opposing gender patterns illustrate that the intergenerational transmission process may vary across generations, which constitutes a challenge for attempts to estimate transmission models from multigenerational data (Nybom and Stuhler, 2019). Figure A.2 in the Appendix shows the corresponding distributions for earnings rank slope coefficients. In that case, the father-child coefficients are larger than those for mother-child pairs, reflecting that mothers’ earnings tend to be poorer proxies of underlying social status. Nevertheless, the differences between the two are quite small, reflecting that the mothers in our sample were already participating on the labor market to a comparatively large degree. Moreover, the grandfather-child coefficients follow a strikingly similar distribution for paternal and maternal-side grandfathers.
@@ 3.1 Comparing inter- and multigenerational mobility
Parent-child correlations as plotted in Figure 1 may understate the extent to which socio-economic advantages are transmitted from one generation to the next, because educational attainment is just one aspect of a person’s overall socio-economic status. Indeed, studies that combine multiple status proxies (Vosters and Nybom, 2017; Blundell and Risa, 2018), that integrate information from grandparents or the extended family (Braun and Stuhler, 2018; Colagrossi, d’Hombres and Schnepf, 2020; Adermon, Lindahl and Palme, 2021; Collado, Ortuño-Ortín and Stuhler, 2023), or that consider long-run mobility across many generations (Clark, 2014; Barone and Mocetti, 2021) tend to find higher persistence.
We illustrate this argument in two ways. First, we use generational averages rather than individual outcomes of a single parent or grandparent. In particular, we compute mean outcomes for both parents and all four grandparents, forming generation-specific means of either schooling or earnings ranks. By averaging over both parents, or all grandparents, these mean measures might provide a better proxy for socioeconomic status of a family. They might also reduce the influence of measurement and sampling error. [^12] Previous studies that construct averages across several family members of the same generation or even across the entire family tree include Thaning and Hällsten (2020), Adermon, Lindahl and Palme (2021), and Hällsten and Kolk (2023).
@@
As already alluded to, the correlations in Table 2 also vary with the gender of the child (see Table A.2 in the Appendix); for example, the (weighted) father-son correlation in years of schooling is on average about 0.33, while the father-daughter correlation is slightly lower at 0.27; and this difference is statistically significant ($p<0.01$). While the pattern flips when considering mothers only, the correlations for sons are slightly larger once the schooling of both parents are taken into account. Overall, the relative differences along gender-specific lines should however be regarded as fairly moderate. Sons also appear slightly less multigenerationally mobile; the schooling correlations tend to be larger for sons irrespectively of whether we consider individual grandparents or jointly take all grandparents into account (0.16 for sons vs. 0.13 for daughters, $p<0.01$). Apart from these differences in levels, there are few gender-specific differences in the multigenerational case. The patterns are largely similar for earnings ranks, with gender-specific transmission in the inter- but less so in the multigenerational case. [^14] Engzell and Mood (2023) and Brandén, Nybom and Vosters (2024) show similar gender-specific patterns for intergenerational persistence in earnings ranks in Sweden. We study below if regional rankings are similar when considering different genders and lineages.
@@
Second, we exploit the availability of three-generation data and compare the relative sizes of our inter- and multigenerational correlations. Recent studies on the national level find that multigenerational correlations tend to be larger than a simple iteration of parent-child correlations would suggest, i.e. intergenerational transmission cannot be approximated by a first-order Markov process; see Anderson, Sheppard and Monden (2018) and Stuhler (2024) for recent surveys. [^15] Moreover, see Lindahl et al. (2015), Dribe and Helgertz (2016), Hällsten and Pfeffer (2017), Engzell, Mood and Jonsson (2020), Adermon, Lindahl and Palme (2021), Helgertz and Dribe (2022), Collado, Ortuño-Ortín and Stuhler (2023) and Hällsten and Kolk (2023) for multigenerational evidence from Sweden.
@@
We further conduct formal statistical tests of $\Delta=0$ across municipalities. [^17] Since $\Delta$ is a non-linear function of parameters, we use the delta rule to derive its approximate standard error and t statistics separately for each municipality, $Var(\triangle)$ $\approx\left(\frac{\partial\triangle}{\partial\beta_{-2}}\right)Var(\beta_{-2})+\left(\frac{\partial\triangle}{\partial\beta_{-1}}\right)^{2}Var(\beta_{-1})+2\left(\frac{\partial\triangle}{\partial\beta_{-2}}\right)\left(\frac{\partial\triangle}{\partial\beta_{-1}}\right)Cov(\beta_{-2},\beta_{-1})$ $\approx Var(\beta_{-2})+(2\beta_{-1})^{2}Var(\beta_{-1})-2(2\beta_{-1})Cov(\beta_{-2},\beta_{-1})$ where we retrieve estimates of $Cov(\beta_{-2},\beta_{-1})$ from the robust covariance matrix generated by seemingly unrelated regressions (SUR). The test statistic is then given by $T=\frac{\triangle}{\sqrt{Var(\triangle)}}$ and can be compared to standard normal critical values. For education, the t statistics are larger than 1.96 in about 33% (unweighted) and 55% (weighted by the number of pairs) of municipalities. For earnings ranks, the corresponding shares are 42% and 67%. Figures A.3a and A.3b in the Appendix plot the full distribution of the test statistics for both education and earnings ranks, showing that $\Delta$ is virtually never significantly negative. Consequently, the implied rejection rates for the one-sided hypothesis $\Delta\leq 0$ are high.

### v2

@@ Abstract
Using complete-count register data spanning three generations, we document spatial patterns in inter- and multigenerational mobility in Sweden. Across municipalities, grandfather-child correlations in education or earnings tend to be larger than the square of the parent-child correlations, suggesting that the latter understate status transmission in the long run. Yet, conventional parent-child correlations capture regional differences in long-run transmission and therefore remain useful for comparative purposes. We further find that the within-country association between mobility and income inequality (the “Great Gatsby Curve”) is at least as strong in the multi- as in the intergenerational case. Interpreting those patterns through the lens of a latent factor model, we find that regional differences in mobility primarily reflect variation in the transmission of latent advantages, rather than in how those advantages translate into observed outcomes.
@@ 1 Introduction
An increasingly popular approach to study intergenerational mobility is to compare mobility rates across regions. Theoretical models with social or spatial dimensions (e.g. Durlauf1996 and Benabou1996a) suggest that income persists across generations partly because parents shape their children’s social environment, such as schools or neighborhoods. Empirically, regional variations can be used to identify causal effects; for example, Holmlund2006 and PekkarinenUusitaloKerr2009 exploit the staggered implementation of an educational reform to estimate its impact on intergenerational mobility. Following CHKS_Where_QJE, some studies invert this research design: rather than focusing on a specific policy or event, they first map regional variation in intergenerational mobility and then search for regional correlates that could explain this variation. [^1] Examples include ConnollyCorakHaeck2019JOLE, ConnollyHaeckLapierre2019 and Corak2017Landscapes for Canada, DeutscherMazumder2020Labour for Australia, Heidrich:2017aa and Branden2019 for Sweden, Risa:2019aa and butikofer2022breaking for Norway, ERIKSEN2020109024 for Denmark, acciari2022and for Italy, or BellBlundellMachin2023 for the UK.
In this study, we extend this “regional approach” by considering not only inter- but also multigenerational associations. Such associations are interesting both from a descriptive and a causal perspective. Recent studies indicate that socioeconomic disparities are more persistent than extrapolations from the available parent-child correlations would suggest (AndersonSheppardMonden2017; Solon_2018; stuhler2024multigenerational). This contrast between inter- and multigenerational correlations is in turn informative about the statistical properties and potential mechanisms of mobility processes (e.g., ColladoOrtunoStuhlerKinship). By adopting such multigenerational perspective, we aim to better understand the causes and implications of regional differences in mobility.
@@
We first report regional estimates of educational and earnings mobility across three generations for cohorts born in the 1980s across 290 municipalities. Municipalities are key geographical entities in the Swedish context, as many public goods and schooling are provided at this level, and often correspond to the relevant labor market of their inhabitants. Our evidence here is complementary to evidence provided by Heidrich:2017aa and Branden2019, who estimate income mobility for earlier birth cohorts and a broader definition of local labor markets or commuting zones.
@@
We then study whether such regional rankings are sensitive to the choice of outcome (education vs. earnings), statistic (e.g., regression vs. correlation coefficients), time frame (inter- vs. multigenerational), or lineage (paternal vs. maternal line). One particular question is whether summary statistics based on education – which can be estimated in most settings – can serve as a substitute for more data-demanding measures based on income. This part of our analysis is complementary to recent work by deutscher2023measuring, who compare the ranking of Australian regions across different measures of income mobility.
@@
We find that regional rankings based on inter- and multigenerational measures are similar. Municipalities with low intergenerational mobility in education (high parent-child correlations) also tend to have low multigenerational correlations (high grandparent-child correlations). In contrast, the association between education- and earnings-based mobility measures is weak. While this observation contrasts with cross-national comparisons, in which educational and income mobility are found to be closely related (Blanden2011Survey; Stuhler2018JRC), it is reminiscent of the sociological discourse on the “mobility paradox”, with some countries showing diverging patterns of income compared to educational or occupational mobility (e.g. breen2016much; karlson2021denmark).
@@
We next study whether multigenerational mobility and cross-sectional inequality are systematically related across regions. A well-documented pattern in international comparisons shows a negative association between income mobility and inequality (Blanden2011Survey, CorakJEP2013). We find that the same relationship holds across municipalities within Sweden, not only for inter- but also for multigenerational mobility. In line with evidence by Branden2019, who analyzed father-child mobility for earlier Swedish cohorts, we observe that municipalities that are characterized by high income inequality in the parent’s generation also tend to exhibit low mobility. However, the strength of this relationship varies considerably with the choice of mobility statistic. Educational mobility is more strongly correlated with inequality than income mobility, possibly due to greater measurement error in income data. Interestingly, the relationship with inequality is particularly pronounced for multigenerational correlations in education.
@@ 2.1 Samples and outcomes
Our measure of labor earnings includes the annual sum (before taxes and transfers) of all earnings according to employer-reported tax and earnings statements to the national tax agency, which captures all types of labor compensation subject to payroll taxes. [^5] We observe labor earnings (including income from self-employment) for all residents in 1968, 1970, 1971, 1973, 1975, 1976, 1979, 1980, 1982 and annually for 1985-2020. We adjust annual incomes for inflation using the CPI. While it would have been interesting to use disposable income, the information on disposable income does not go as far back in time. The measure is not censored or top-coded, and includes bonus payments. For our purposes, it is crucial to use earnings measures that are as comparable as possible over generations, as our estimates would be distorted if the influence of lifecycle and attenuation biases varies across generations. Given the structure of our data set, with earnings observed over a given time period and different generations (naturally) born several decades apart, it is not possible to measure earnings over the same age ranges for all generations. We therefore use a Mincer-type approach to predict prime-age earnings for all generations, which accounts for differences in income growth across education groups and differences in income levels across cohorts. [^6] While this approach improves our earnings measures, they are still unlikely to be fully comparable across generations. Specifically, we do not account for the observation that age-education-earnings profiles may differ across cohorts (heckman2003fifty) or be steeper for children from parents with higher socioeconomic status (MelloNybomStuhlerLifecycle). However, we do allow for individual intercepts, thereby capturing general shifts in earnings levels across generations. Table A.1 in the Appendix compares area-level inter- and multigenerational earnings rank slopes for our prediction-based measure with more standard measures based on observed incomes. While the dispersion across municipalities is largely similar across measures, the prediction-based measure yields larger coefficients with magnitudes that are more in line with recent evidence for Sweden.
@@ 2.2 Mobility statistics
where $y_{ir}$ is one of the outcomes described above for individual $i$ in region (municipality) $r$ and $y_{i}^{p}$ is the corresponding outcome for a parent or grandparent. The slope $\beta_{r}$ measures “relative” mobility, comparing socioeconomic status relative to the status of others in the same generation. For education, we also consider the corresponding Pearson correlation coefficient, which abstracts from shifts in the variance between the two generations within the region. For our analysis of earnings mobility, we consider either log earnings (in which case the slope coefficient $\beta_{r}$ can be interpreted as the intergenerational elasticity, i.e. the IGE) or earnings ranks (in which case $\beta_{r}$ can be referred to as the rank slope). Ranks are defined on the national level, separately within each child cohort. [^10] Based on the estimated intercept $\hat{\alpha}_{r}$ and the slope $\hat{\beta}_{r}$ we further construct the expected “absolute upward mobility” of children born to parents at the 25th percentile (i.e., for $y_{i}^{p}=0.25$), as defined in CHKS_Where_QJE. Our primary focus is on education correlations or the earnings rank slope, and we focus only occasionally on the IGE and the measure of upward mobility.
@@ 3 Geographical Variation in Mobility
As daughters’ education resembles their mother’s more than their father’s education (the relations are more symmetric for sons), mother-child correlations tend to be marginally larger than the corresponding father-child correlations (0.31 vs 0.30); see Appendix Table A.2 for separate statistics by child gender. The inverse pattern applies for grandparent correlations, which are systematically larger for grandfather-child than for grandmother-child pairs (0.12 vs 0.09). [^11] These opposing gender patterns illustrate that the intergenerational transmission process may vary across generations, which constitutes a challenge for attempts to estimate transmission models from multigenerational data (NybomStuhler2019). Figure A.2 in the Appendix shows the corresponding distributions for earnings rank slope coefficients. In that case, the father-child coefficients are larger than those for mother-child pairs, reflecting that mothers’ earnings tend to be poorer proxies of underlying social status. Nevertheless, the differences between the two are quite small, reflecting that the mothers in our sample were already participating on the labor market to a comparatively large degree. Moreover, the grandfather-child coefficients follow a strikingly similar distribution for paternal and maternal-side grandfathers.
@@ 3.1 Comparing inter- and multigenerational mobility
Parent-child correlations as plotted in Figure 1 may understate the extent to which socio-economic advantages are transmitted from one generation to the next, because educational attainment is just one aspect of a person’s overall socio-economic status. Indeed, studies that combine multiple status proxies (VostersNybom2017; BlundellRisa2018aa), that integrate information from grandparents or the extended family (ECOJ:ECOJ12453; colagrossi2020like; Adermonetal2016; ColladoOrtunoStuhlerKinship), or that consider long-run mobility across many generations (Clark2014book; BaroneMocetti2016) tend to find higher persistence.
We illustrate this argument in two ways. First, we use generational averages rather than individual outcomes of a single parent or grandparent. In particular, we compute mean outcomes for both parents and all four grandparents, forming generation-specific means of either schooling or earnings ranks. By averaging over both parents, or all grandparents, these mean measures might provide a better proxy for socioeconomic status of a family. They might also reduce the influence of measurement and sampling error. [^12] Previous studies that construct averages across several family members of the same generation or even across the entire family tree include thaning2020end, Adermonetal2016, and hallsten2023shadow.
@@
As already alluded to, the correlations in Table 2 also vary with the gender of the child (see Table A.2 in the Appendix); for example, the (weighted) father-son correlation in years of schooling is on average about 0.33, while the father-daughter correlation is slightly lower at 0.27; and this difference is statistically significant ($p<0.01$). While the pattern flips when considering mothers only, the correlations for sons are slightly larger once the schooling of both parents are taken into account. Overall, the relative differences along gender-specific lines should however be regarded as fairly moderate. Sons also appear slightly less multigenerationally mobile; the schooling correlations tend to be larger for sons irrespectively of whether we consider individual grandparents or jointly take all grandparents into account (0.16 for sons vs. 0.13 for daughters, $p<0.01$). Apart from these differences in levels, there are few gender-specific differences in the multigenerational case. The patterns are largely similar for earnings ranks, with gender-specific transmission in the inter- but less so in the multigenerational case.141414engzell2023understanding and branden2024like show similar gender-specific patterns for intergenerational persistence in earnings ranks in Sweden. We study below if regional rankings are similar when considering different genders and lineages.
@@
Second, we exploit the availability of three-generation data and compare the relative sizes of our inter- and multigenerational correlations. Recent studies on the national level find that multigenerational correlations tend to be larger than a simple iteration of parent-child correlations would suggest, i.e. intergenerational transmission cannot be approximated by a first-order Markov process; see AndersonSheppardMonden2017 and stuhler2024multigenerational for recent surveys. [^15] Moreover, see LindahlPalme2014_IGE4Generations, Dribe:2016aa, hallsten2017grand, Engzell:2020aa, Adermonetal2016, HelgertzDribe2021, ColladoOrtunoStuhlerKinship and hallsten2023shadow for multigenerational evidence from Sweden.
@@
We further conduct formal statistical tests of $\Delta=0$ across municipalities. [^17] Since $\Delta$ is a non-linear function of parameters, we use the delta rule to derive its approximate standard error and t statistics separately for each municipality, $Var(\triangle)$ $\approx\left(\frac{\partial\triangle}{\partial\beta_{-2}}\right)Var(\beta_{-2})+\left(\frac{\partial\triangle}{\partial\beta_{-1}}\right)^{2}Var(\beta_{-1})+2\left(\frac{\partial\triangle}{\partial\beta_{-2}}\right)\left(\frac{\partial\triangle}{\partial\beta_{-1}}\right)Cov(\beta_{-2},\beta_{-1})$ $\approx Var(\beta_{-2})+(2\beta_{-1})^{2}Var(\beta_{-1})-2(2\beta_{-1})Cov(\beta_{-2},\beta_{-1})$ where we retrieve estimates of $Cov(\beta_{-2},\beta_{-1})$ from the robust covariance matrix generated by seemingly unrelated regressions (SUR). The test statistic is then given by $T=\frac{\triangle}{\sqrt{Var(\triangle)}}$ and can be compared to standard normal critical values. For education, the t statistics are larger than 1.96 in about 33% (unweighted) and 55% (weighted by the number of pairs) of municipalities. For earnings ranks, the corresponding shares are 42% and 67%. Figure A.3 in the Appendix plots the full distribution of the test statistics for both education and earnings ranks, showing that $\Delta$ is virtually never significantly negative. Consequently, the implied rejection rates for the one-sided hypothesis $\Delta\leq 0$ are high.